OpenStudy (anonymous):

find th maximum y-value on the graph on y=f(x). f(x)=-x^2+10x+3 the maximum y-value is y=

OpenStudy (anonymous):

First, find derivative

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

What course are you taking? what mrtods have you been using?

OpenStudy (anonymous):

Algebra 2

OpenStudy (anonymous):

are you able to help me

OpenStudy (anonymous):

First find the x coordinate of the vertex\[\frac{ b }{ 2a }\]

OpenStudy (anonymous):

how do I do that

OpenStudy (anonymous):

\[ax ^{2}+bx+c=0\]

OpenStudy (anonymous):

okwhat do I do now

OpenStudy (anonymous):

\[-x^{2}+10x+3\]\[ax ^{2}+bx+c\]a=-1 b=10 c=3 Do you take notes in class?

OpenStudy (anonymous):

yes I do take notes but having problems undertanding it, so what woud y=

OpenStudy (anonymous):

So you find x-coodinate of the vertex\[\frac{ b }{ 2a }\]

OpenStudy (anonymous):

10*2

OpenStudy (anonymous):

Take it one step at a time. Just plug in the values in \[\frac{ b }{ 2a }\]

OpenStudy (anonymous):

do not understand

OpenStudy (anonymous):

Just read above, I told you what is a, what is b. Just put the values of a and b in\[\frac{ b }{ 2a }\]

OpenStudy (anonymous):

b is 10 anda s 1

OpenStudy (anonymous):

You put it in the expression\[\frac{ 10 }{ 2(-1) }\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what do i do nxt

OpenStudy (anonymous):

That expression simplifies to a number. You need to get that number.

OpenStudy (anonymous):

can you tell mehow to simplifies te exprsson

OpenStudy (anonymous):

would it be 5

OpenStudy (anonymous):

You need one-on-one time with your teacher or a tutor. Print out these messages and tell them you are having trouble following.

OpenStudy (anonymous):

wtever guess you cannot hep me

OpenStudy (anonymous):

I did cotact a tutor and they s nt helpful